OpenStudy (dls):

Find nth derivative of tan^(-1) (x/a).

OpenStudy (dls):

@hartnn

OpenStudy (dls):

Can't observe any visible pattern from my side. http://www.wolframalpha.com/input/?i=Table%5BD%5BArcTan%5Bx%2Fa%5D%2C+%7Bx%2C+n%7D%5D%2C+%7Bn%2C+1%2C+10%7D%5D

OpenStudy (dls):

@rational

OpenStudy (dls):

@nincompoop

OpenStudy (dls):

what on earth

OpenStudy (rational):

pick the one that looks easy, we can anayze..

OpenStudy (dls):

none :P

OpenStudy (rational):

same here lol

OpenStudy (rational):

looks wolfram is interpreting x as complex number

OpenStudy (dls):

hmm :| anything better?

OpenStudy (rational):

let me grab pen and paper, i feel its going to be another 3-4 pages >.<

OpenStudy (dls):

hehe XD

OpenStudy (rational):

\[\large \begin{align}\\ y &= \arctan(x/a) \\~\\ y_1/a &= \dfrac{1}{a^2+x^2} = \dfrac{0!}{a^2+x^2} \\~\\ y_2/a &= \dfrac{-2x}{(a^2+x^2)^2} = -\dfrac{1!. \binom{2}{1}x}{a^2+x^2} \\~\\ y_3/a &= \dfrac{2(3x^2-a^2)}{(a^2+x^2)^3} = \dfrac{2!.\left( \binom{3}{1}x^2-\binom{3}{3}a^2\right)}{(a^2+x^2)^3} \\~\\ y_4/a &= -\dfrac{24(x^3-a^2x)}{(a^2+x^2)^4} = \dfrac{3!.\left( \binom{4}{1}x^3-\binom{4}{3}a^2x\right)}{(a^2+x^2)^4} \\~\\ \cdots\\~\\ y_n/a &= \dfrac{(-1)^n(n-1)!\left(\binom{n}{1}x^{n-1}a^0 -\binom{n}{3}x^{n-3}a^2 + \cdots \right)}{(a^2+x^2)^n} \end{align}\]

OpenStudy (rational):

Im not going to lie, I have just reverse calculated it from wolfram :) you're an einstein if you make sense of above stuff ;)

OpenStudy (dls):

lets let it be :P